IDENTIFICATION OF PRIMARY AND SECONDARY PHOTOPRODUCTS OF ULTRAVIOLET PHO-TODISSOCIATION OF 1,3-BUTADIENE

\underline{JAIME A. STEARNS}, ALOKE DAS, AND TIMOTHY S. ZWIER, Department of Chemistry, Purdue University, West Lafayette, IN 47907.

The recombination of propargyl radicals is thought to be an important route to the formation of benzene both in planetary atmospheres and in flames. 1,3-Butadiene is one likely source of propargyl radical in such environments. We have carried out a study of the primary and secondary products of photodissociation of 1,3-butadiene following excitation near 220 nm in a supersonic expansion. Photoexcitation occurred inside a one-centimeter long reaction tube affixed to the nozzle of the pulsed valve, providing a 20μs window during which primary dissociation products could undergo recombination before expanding. The major primary products in the 118 nm ionization time-of-flight mass spectrum were C_{3}H_{3}, C_{4}H_{4}, and C_{4}H_{5}. The recombination of two C_{3}H_{3} radicals, giving C_{6}H_{6}, was shown to be the major secondary reaction channel. Other secondary products were C_{5}H_{8}, C_{7}H_{7}, and C_{7}H_{8}. The primary and secondary products were characterized by resonant two-photon ionization and resonant ion-dip infrared spectroscopies.